### v1

# Decoupling Strategy and Execution in Task-Focused Dialogue via Goal-Oriented Preference Optimization

Abstract
Large language models show potential in task-oriented dialogue systems, yet existing training methods often rely on token-level likelihood or preference optimization, which poorly align with long-horizon task success. To address this, we propose Goal-Oriented Preference Optimization (GOPO), a hierarchical reinforcement learning framework that decouples strategy planning from response generation via an Expert Agent and a Customer Service Agent. The Expert Agent optimizes multi-turn goal preferences at the dialogue-trajectory level, while the Customer Service Agent generates responses strictly aligned with the selected strategy. We evaluate GOPO on public benchmarks and e-commerce customer service datasets, and introduce Task-focused Sequential Engagement (TSE), a sequence-level metric derived from real e-commerce interaction data. On the Mgshop dataset, GOPO improves TSE by 7.7% and 10.3% over PPO and Memento, with consistent gains in sequence-level reward and generation quality. Furthermore, a 14B model trained with GOPO achieves 2.7% and 1.5% higher TSE than Qwen-235B and GPT-5.2, respectively. Ablation studies confirm the Expert Agent’s critical role in long-horizon optimization. GOPO demonstrates consistent improvements across other datasets as well. This work establishes a new paradigm for task-oriented dialogue systems in commercial scenarios, with code and datasets to be made public.
Machine Learning, ICML
1 Introduction
Recent advances in large language models have substantially improved task-oriented dialogue systems in fluency and language understanding(Brown et al., 2020). However, in business-critical scenarios such as e-commerce and customer service, these systems often fall into an “efficiency trap”: despite strong technical metrics, they yield limited gains on core business KPIs, including conversion rate, user satisfaction, and first-contact resolution. This gap between linguistic proficiency and business effectiveness exposes fundamental limitations of current research paradigms(Touvron et al., 2023).
Figure 1: TSE and Reward Across Datasets: GOPO-Qwen3-14B vs. Qwen-235B, DeepSeek-R1, and GPT-5.2.
We identify three interrelated challenges underlying this issue. First, optimization objectives are misaligned with business goals. Mainstream approaches such as supervised fine-tuning and RLHF prioritize technical criteria like preference alignment and semantic relevance, while overlooking business outcomes, resulting in fluent yet ineffective responses that fail to drive conversion or resolve complex issues. Second, existing single-agent architectures lack strategic flexibility in complex multi-turn dialogues. By conflating strategy selection with response generation, they struggle to adapt to dynamic user intentions and emotional shifts, leading to repetitive or fragmented interactions and degraded user experience. Finally, standard operating procedure (SOP) compliance remains unreliable. Prompt-based soft constraints cannot consistently enforce enterprise rules such as mandatory disclaimers or commitment prohibitions, causing strategic intent to be diluted during execution and leading to unstable service quality and compliance risks.
To address these challenges, we propose GOPO, a hierarchical reinforcement learning (RL) framework for task-oriented dialogue. GOPO adopts a dual-agent architecture that decouples strategy planning from response execution, comprising an Expert Agent for high-level decision-making and a Customer Service Agent for compliant response generation. The Expert Agent analyzes dialogue states and dynamically assembles strategy chains from a skill pool, which are translated into structured hard constraints and directly enforced by the Customer Service Agent, ensuring strict alignment between strategic intent and generated responses. Moreover, GOPO employs a joint reward model tied to business end-state metrics, aligning long-horizon optimization with real business value.
The primary contributions of this work are summarized as follows:
  • We propose GOPO, a hierarchical RL framework that decouples strategy planning from response generation via an Expert Agent and a Customer Service Agent, enabling long-horizon optimization in task-oriented dialogues.
  • We introduce Task-focused Sequential Engagement (TSE), a sequence-level evaluation metric derived from real e-commerce interactions, designed to better reflect long-horizon task success.
  • Extensive experiments on public benchmarks and real-world e-commerce customer service datasets show that GOPO consistently outperforms PPO, Memento, and strong large-model baselines in sequence-level reward, generation quality, and TSE.
  • Results further demonstrate that GOPO enables smaller models to outperform significantly larger models, and ablation studies confirm the Expert Agent’s key role in long-horizon preference optimization.
The remainder of this paper is organized as follows. Section 2 reviews the related work. Section 3 presents the methodology of the GOPO framework in detail. Section 4 presents the experimental setup, results, and an in-depth analysis of the experimental findings. Finally, Section 5 concludes the paper and outlines directions for future research.
2 Related works
2.1 Agent Architectures in Task-Oriented Dialogue Systems
Task-oriented dialogue systems in e-commerce are typical carriers of business value orientation(Sanjaya et al., 2023; Konda et al., 2021; Gamboa-Cruzado et al., 2023). Their core goal is to achieve a closed loop of demand matching, trust building and order conversion via multi-turn interactions, making this a major research focus recently(Kanojiya et al., 2021). Existing studies on e-commerce dialogue systems fall broadly into two categories: traditional rule-based or modular pipeline methods, and LLM-based end-to-end single-agent approaches(Puertas et al., 2024). Traditional systems adopt a modular pipeline architecture, sequentially executing natural language understanding (NLU), dialogue management (DM) and natural language generation (NLG). With the rise of LLMs, however, end-to-end single-agent architectures have become dominant. Represented by ReAct(Wei et al., 2025) and its variants, these architectures use LLMs as central controllers and implement a “think–act” loop for complex reasoning and tool-use tasks.
Nevertheless, these systems face limitations in complex, long-horizon business processes. A single agent must both select SOP-driven strategies and generate compliant responses, tightly coupling decision logic with execution constraints. Under uncertainties such as ambiguous user intent or emotional fluctuations, this leads to suboptimal strategies, compliance failures, or dialogue stagnation.
To overcome single-agent limitations, the community has explored multi-agent approaches. CAMEL(Li et al., 2023) uses a role-playing framework where two agents debate solutions, while other works design agent teams with explicit role divisions, e.g., a router agent dispatching requests. Although these methods partially decompose tasks, they rely on natural language soft constraints for inter-agent communication, which is unreliable in strict SOP scenarios, as LLM stochasticity can dilute strategic intent. In contrast, GOPO employs a hierarchical dual-agent architecture with structured hard-constraint transmission, ensuring Expert Agent’s strategies are executed by the Customer Service Agent.
2.2 Optimization Methods for Dialogue Policies
RL has long been a core approach for optimizing dialogue policies (Peiyuan et al., 2024). Classical methods such as Proximal Policy Optimization (PPO) maximize cumulative rewards, but in multi-turn dialogues rewards are sparse and delayed, with transaction success only observable at the end of a conversation. This leads to severe long-horizon credit assignment issues, making optimization inefficient and prone to local optima, such as over-emphasizing fluency while neglecting SOP compliance.
To mitigate reward design challenges, preference-based optimization methods, particularly Direct Preference Optimization (DPO) and its variants, have gained attention. DPO leverages pairwise human preferences to avoid explicit reward modeling and online RL, and subsequent frameworks (Wu et al., 2025) extend this paradigm to real-time dialogue settings using large-scale offline data. Building on this line of work, Single-stream Policy Optimization (SPO) (Xu and Ding, 2025) and Stepwise Direct Preference Optimization (SDPO) (Kim et al., 2025) introduce finer-grained supervision. However, these approaches rely on static preferences and primarily optimize technical alignment metrics, without explicitly incorporating outcome-driven business signals such as user satisfaction or purchase conversion. Recent methods such as (Zhou et al., 2025) further explore trajectory-level optimization but remain focused on generic preference consistency rather than task-specific business objectives. In contrast, GOPO adopts a joint reward model directly tied to business metrics and, through a hierarchical RL design, provides timely signals that effectively address long-horizon credit assignment.
2.3 SOP Compliance and Constraint Enforcement
In heavily regulated industries such as finance, healthcare and e-commerce, strict adherence to standard operating procedures is critical for dialogue systems. Existing methods enforce constraints mainly via prompt engineering or rule-based post-processing filtering. However, prompt-based ”soft constraints” are fragile against LLMs’ generative creativity and cannot ensure full compliance [16]. Post-processing filters are passive remedial measures that fail to fundamentally prevent non-compliant outputs and may impair response fluency and naturalness.
Some studies attempt to introduce stronger constraints during decoding. For example, constrained beam search(Chen and Authors, 2025) or vocabulary masking during decoding can enforce the generation of specific keywords or prevent prohibited terms. While these methods improve compliance to some extent, they generally struggle with complex logical constraints and may conflict with the model’s internal knowledge, leading to degraded generation quality.
Our GOPO framework addresses this issue through a novel hard-constraint transmission mechanism. The Expert Agent not only decides the SOP strategy but also compiles it into a set of structured hard constraints, which are directly integrated into the Customer Service Agent’s generation process. Additionally, we design an SOP matching reward module that provides real-time validation and rewards for compliance at the execution layer, creating a complete closed loop from strategy decision-making to response generation and ensuring strict adherence to SOPs.
Figure 2: Hierarchical Dual-Agent Architecture of GOPO with Decoupled Strategy and Generation.
3 The GOPO Framework
In this section, we introduce the GOPO framework, formalize the task-oriented dialogue problem under business constraints, and present a joint reward and loss design that aligns hierarchical RL with practical business outcomes.
3.1 Framework Overview
Figure 2 presents the hierarchical dual-agent architecture of GOPO. The architecture consists of an Expert Agent for strategy planning and a Customer Service Agent for constrained response generation. Under a HRL paradigm, the Expert Agent composes skills from a predefined pool, decoupling policy planning from content generation.
  1. 1.
    State Perception: The system extracts dialogue state information (e.g., user intent and emotion).
  2. 2.
    Skill Selection: Given the global state $s_{t}^{E}$, $E$ selects a skill sequence as a macro-action $a_{t}^{E}$ to define the turn-level strategy.
  3. 3.
    Response Generation: Conditioned on $a_{t}^{E}$ and local state $s_{t}^{A}$, $A$ generates a compliant response $a_{t}^{A}$ under hard constraints.
  4. 4.
    Feedback and Transition: The environment returns a joint reward $R_{t}$ and transitions to $s_{t+1}$.
This hierarchical division decomposes dialogue tasks into two coupled subproblems, allowing independent optimization of strategy and response generation.
3.2 Problem Formalization
We formalize the dynamic SOP task-oriented dialogue process as a Hierarchical Markov Decision Process (HMDP), represented by the tuple
$\text{HMDP}=\langle S_{E},S_{A},A_{E},A_{A},R_{E},R_{A},R_{\text{total}}\rangle,$
where independent MDPs are defined for $E$ and $A$, coordinated via hierarchical dependencies between states and actions. Each element is defined as follows:
  • State of Expert Agent ($s_{t}^{E}\in S_{E}$): Captures global and temporal dialogue information. At time $t$, the state $s_{t}^{E}$ includes:
    • $H_{t}$: Dialogue history of the most recent $k$ turns.
    • $u_{t}$: Current user intent label.
    • $e_{t}$: Current user emotion label.
    • $a_{t-1}^{E}$: The skill selected by $E$ to establish temporal dependency in decision-making.
  • Action ($a_{t}^{E}\in A_{E}$): The action space of $E$. $A^{E}$ is discrete, representing the selection of one or more skills from the skill pool $P_{\text{skill}}$. Therefore, $a_{t}^{E}\in P_{\text{skill}}$.
  • Reward ($R_{t}^{E}\in R_{E}$): The reward for $E$. $R_{t}^{E}$ evaluates the long-term value of its skill choices.
  • State of Customer Service Agent($s_{t}^{A}\in S_{A}$): $S_{A}$ focuses on execution details of the current turn. $s_{t}^{A}$ includes:
    • $U_{t}$: The user’s utterance in the current turn.
    • $a_{t}^{E}$: The skill selected by $E$, serving as a hard execution constraint.
    • $B_{t}$: Relevant business context information, such as order status or inventory.
  • Action ($a_{t}^{A}\in A_{A}$): The action space of $A$. $A^{A}$ is continuous, representing the generated response text sequence. $a_{t}^{A}=(w_{1},w_{2},\dots,w_{m})$, where $w_{i}$ is a token from the vocabulary.
  • Reward ($R_{t}^{A}\in R_{A}$): The reward for $A$, used to evaluate the quality and compliance of the response.
  • Joint Reward ($R_{t}\in R_{\text{total}}$): A unified reward that jointly evaluates the strategy decision of $E$ and the response execution of $A$. This reward signal is propagated back to both agents. Detailed discussion is provided in Section 3.3.
Based on the MDP formulation, GOPO aims to select skills to define turn-level strategies and generate constrained responses that satisfy immediate dialogue needs while maximizing long-term business objectives under strict compliance.
Table 1: Comparison of GOPO (Qwen-7B-Chat) with Large-Scale Baselines Across Multiple Datasets (TSE, Reward, BLEU).
Model Mgshop Multiwoz TmallBrand-A TmallBrand-B
TSE Reward BLEU TSE Reward BLEU TSE Reward BLEU TSE Reward BLEU
Untrained 0.7454 5.97 0.091 0.6127 4.53 0.075 0.6791 5.25 0.083 0.6685 5.31 0.082
SFT 0.8363 6.25 0.187 0.7015 4.92 0.149 0.7689 5.59 0.165 0.7592 5.64 0.163
PPO 0.8584 7.09 0.190 0.7496 5.68 0.169 0.8040 6.39 0.181 0.8126 6.43 0.183
Memento 0.8381 7.13 0.188 0.7205 5.75 0.153 0.7793 6.44 0.168 0.7848 6.38 0.169
GOPO(Ours) 0.9243 7.38 0.211 0.7543 6.38 0.172 0.8393 6.88 0.192 0.8417 6.92 0.193
Table 2: Comparison of GOPO with Large-Scale Baselines Across Multiple Datasets (TSE, Reward, BLEU).
Model Mgshop Multiwoz TmallBrand-A TmallBrand-B
TSE Reward BLEU TSE Reward BLEU TSE Reward BLEU TSE Reward BLEU
Qwen-235B 0.9227 7.24 0.182 0.7462 5.57 0.165 0.8327 6.38 0.114 0.8294 6.42 0.166
DeepSeek-R1 0.9381 7.46 0.143 0.7609 6.46 0.138 0.8512 7.44 0.141 0.8489 7.41 0.139
GLM-4.7 0.9365 7.25 0.150 0.7598 6.13 0.148 0.8487 7.09 0.153 0.8301 6.36 0.151
GPT-5.2 0.9338 7.54 0.097 0.7561 5.83 0.092 0.8426 6.72 0.095 0.8453 6.71 0.094
Gemini-2.5 0.9287 7.35 0.133 0.7451 6.16 0.129 0.8084 6.91 0.131 0.8116 6.88 0.128
GOPO-Qwen3-14B(Ours) 0.9475 7.63 0.279 0.7699 6.55 0.177 0.8511 7.46 0.186 0.8524 7.43 0.198
Accordingly, GOPO adopts a hierarchical decision mechanism in which $E$ learns a macro-policy $\pi_{E}(a_{t}^{E}\mid s_{t}^{E})$ to select skill combinations that guide the dialogue toward business goals, while $A$ learns a micro-policy $\pi_{A}(a_{t}^{A}\mid s_{t}^{A})$ to generate high-quality, compliant responses that fit the current dialogue context. Formally, the optimization objective of GOPO is to train the joint policy $\pi=\langle\pi_{E},\pi_{A}\rangle$ to maximize the expected cumulative reward of business outcome metrics, expressed as:
$\max_{\pi_{E},\pi_{A}}\;\mathbb{E}_{\tau\sim\pi}\left[\sum_{t=0}^{T}\gamma^{t}R_{\text{total}}(s_{t},a_{t}^{E},a_{t}^{A})\right].$ (1)
3.3 Reward and Loss Design
The success of the GOPO framework hinges on the design of its reward and loss functions, which closely align the optimization process with actual business objectives.
The Expert Agent reward $R_{t}^{E}$ evaluates skill selection quality via Expert Skill Normalized Discounted Cumulative Gain (ESNDCG), which measures the relevance between the predicted skill sequence $P=[p_{1},\dots,p_{m}]$ and the Teacher-generated reference sequence $T=[t_{1},\dots,t_{n}]$. For any predicted skill $s$, its relevance is defined as:
$\text{Rel}(s)=\begin{cases}n-\text{index}(s\text{ in }T)+1,&\text{if }s\in T,\\ 0,&\text{if }s\notin T.\end{cases}$ (2)
Based on this relevance definition, we compute the Discounted Cumulative Gain (DCG) for the predicted skill sequence $P=[p_{1},p_{2},\dots,p_{m}]$ as:
$\text{DCG}(P)=\sum_{i=1}^{m}\frac{2^{\text{Rel}(p_{i})}-1}{\log_{2}(i+1)}.$ (3)
The Ideal DCG (IDCG) is defined as the DCG of the Teacher sequence itself, which can be expressed:
$\text{IDCG}(T)=\sum_{j=1}^{n}\frac{2^{\text{Rel}(t_{j})}-1}{\log_{2}(j+1)},$ (4)
which represents the theoretical maximum score. The normalized metric, Expert Skill Normalized Discounted Cumulative Gain (ESNDCG)(Zhang et al., 2022), is then computed as
$\text{ESNDCG}(P,T)=\frac{\text{DCG}(P)}{\text{IDCG}(T)}.$ (5)
Finally, the reward for $E$ is defined as:
$R_{t}^{E}=\text{ESNDCG}(P,T).$ (6)
The loss function for the Expert Agent adopts the standard policy gradient form with an added entropy regularization term to encourage exploration:
$\begin{split}\mathcal{L}_{t}(\theta_{E})=&-\mathbb{E}_{t}\big[\hat{A}_{t}^{E}\cdot\log\pi_{E}(a_{t}^{E}\mid s_{t}^{E};\theta_{E})\big]\\ &-\alpha\cdot\mathcal{H}(\pi_{E}(\cdot\mid s_{t}^{E};\theta_{E})).\end{split}$ (7)
Here, $\hat{A}_{t}^{E}$ denotes the advantage function, which is computed using the Actor-Critic architecture.
The Customer Service Agent reward $R_{t}^{A}$ evaluates response quality, compliance, and diversity. We employ a large language model (GPT-4) as an automatic evaluator, which scores each response along multiple predefined dimensions. The final reward is computed as a weighted sum of these dimension scores:
$R_{t}^{A}=\alpha_{1}S_{1}+\alpha_{2}S_{2}+\alpha_{3}S_{3}+\alpha_{4}S_{4},$ (8)
where $\alpha_{i}$ are normalized weights controlling the contribution of each evaluation dimension.
The loss function for the Customer Service Agent is a composite loss, combining policy gradient, skill compliance, and diversity:
$\mathcal{L}_{t}(\theta_{A})$ $=\lambda_{\text{p}}\,\mathcal{L}_{t,\text{p}}(\theta_{A})+\lambda_{\text{s}}\,\mathcal{L}_{t,\text{s}}(\theta_{A})+\lambda_{\text{d}}\,\mathcal{L}_{t,\text{d}}(\theta_{A}),$ (9a)
$\mathcal{L}_{t,\text{p}}(\theta_{A})$ $=-\mathbb{E}_{a_{t}^{A}\sim\pi_{A}}\big[R_{t}^{A}\,\log\pi_{A}(a_{t}^{A}\mid s_{t}^{A})\big],$ (9b)
$\mathcal{L}_{t,\text{s}}(\theta_{A})$ $=d(a_{t}^{A},a_{t}^{E}),$ (9c)
$\mathcal{L}_{t,\text{d}}(\theta_{A})$ $=-\sum_{t}\sum_{w\in V}p_{\theta}(w\mid s_{t}^{A})\,\log p_{\theta}(w\mid s_{t}^{A}),$ (9d)
Here, $\mathcal{L}_{t,\text{p}}$ is the policy gradient loss that maximizes the Customer Service Agent’s expected reward $R_{t}^{A}$, $\mathcal{L}_{t,\text{s}}$ enforces skill compliance via the semantic distance $d(a_{t}^{A},a_{t}^{E})$, and $\mathcal{L}_{t,\text{d}}$ is a negative-entropy term promoting response diversity. The total loss $\mathcal{L}_{t}(\theta_{A})$ is their weighted sum.
Table 3: Comparison of Flagship Models Across Datasets on G-Eval Scores.
Model Mgshop Multiwoz TmallBrand-A TmallBrand-B
Qwen-235B 3.564 3.353 3.753 3.747
DeepSeek-R1 3.476 3.368 3.728 3.745
GLM-4.7 3.672 3.381 3.525 3.510
GPT-5.2 3.529 3.453 3.596 3.517
Gemini-2.5 3.680 3.407 3.653 3.534
GOPO-Qwen3-14B(Ours) 3.688 3.447 3.757 3.761
Table 4: Comparison of GOPO Variants on Multiple Datasets (TSE, Reward, BLEU).
Model Mgshop Multiwoz TmallBrand-A TmallBrand-B
TSE Reward BLEU TSE Reward BLEU TSE Reward BLEU TSE Reward BLEU
GOPO-Qwen-7B-Chat 0.9243 7.38 0.211 0.7543 6.38 0.172 0.8393 6.88 0.192 0.8417 6.92 0.193
GOPO-Qwen3-8B 0.9375 7.44 0.242 0.7680 6.51 0.175 0.8507 7.21 0.202 0.8515 7.43 0.208
GOPO-Qwen3-14B 0.9475 7.63 0.279 0.7699 6.55 0.177 0.8511 7.46 0.186 0.8524 7.43 0.198
To account for shifting dialogue priorities, GOPO uses a dynamic, temporally-aware weighting to combine the Expert and Customer Service Agent rewards:
$R_{t}=w_{t}^{E}\cdot R_{t}^{E}+w_{t}^{A}\cdot R_{t}^{A}.$ (10)
Here, the Expert Agent weight $w_{t}^{E}$ increases over turns, shifting emphasis from early response quality to later strategic skill selection for business goals.
Figure 3: Convergence Performance of GOPO Under Different Parameter Settings.
4 Experiments
To comprehensively evaluate the effectiveness of the GOPO framework, we conducted a series of extensive experiments. This section describes the experimental setup, main results, ablations, and qualitative analysis.
4.1 Experimental Setup
Table 5: Ablation Study of GOPO-14B on Mgshop and TmallBrand-A Datasets.
Model Mgshop TmallBrand-A
TSE Reward BLEU TSE Reward BLEU
Full GOPO 0.9475 7.63 0.279 0.8511 7.46 0.186
GOPO w/o Expert Agent 0.9371 7.16 0.217 0.8023 6.47 0.227
Full Baseline 0.7353 6.92 0.251 0.6371 6.43 0.213
Baseline w/o Expert Agent 0.4327 6.08 0.240 0.1818 5.19 0.182
Table 6: Performance of Customer Service Agents Evaluated with Different Expert Agents on Mgshop Dataset.
Expert Agent/ Customer Service Agent GOPO-Qwen3-14B Qwen-235B DeepSeek-R1 Gemini-2.5
TSE Reward BLEU TSE Reward BLEU TSE Reward BLEU TSE Reward BLEU
GOPO-Qwen3-14B 0.9475 7.63 0.279 0.8714 7.40 0.193 0.7854 6.73 0.234 0.7384 5.30 0.251
Qwen-235B 0.9352 7.45 0.214 0.9239 7.37 0.186 - - - - - -
DeepSeek-R1 0.9453 7.55 0.242 - - - 0.9426 7.51 0.145 - - -
Gemini-2.5 0.9268 7.52 0.220 - - - - - - 0.9265 7.12 0.130
Table 7: Performance of Customer Service Agents Evaluated with Different Expert Agents on TmallBrand-A Dataset.
Expert Agent/ Customer Service Agent GOPO-Qwen3-14B Qwen-235B DeepSeek-R1 Gemini-2.5
TSE Reward BLEU TSE Reward BLEU TSE Reward BLEU TSE Reward BLEU
GOPO-Qwen3-14B 0.8511 7.46 0.186 0.8291 6.27 0.203 0.8395 6.84 0.146 0.8163 6.55 0.228
Qwen-235B 0.8428 7.37 0.201 0.8371 6.54 0.235 - - - - - -
DeepSeek-R1 0.8503 7.03 0.245 - - - 0.8443 7.50 0.218 - - -
Gemini-2.5 0.8025 6.83 0.131 - - - - - - 0.8152 6.87 0.196
4.1.1 Datasets
We evaluate our method on four datasets spanning public benchmarks and real-world enterprise data.
  • Mgshop: An internal test set designed to evaluate robustness under highly dynamic user emotions and rapidly changing intents.
  • Multiwoz(Budzianowski et al., 2018): A widely used multi-domain task-oriented dialogue benchmark for assessing general dialogue capabilities.
  • TmallBrand-A: An internal e-commerce customer service dataset reflecting complex interactions such as purchasing, inquiries, and refunds.
  • TmallBrand-B: An internal e-commerce customer service dataset from a different store, used to assess cross-domain generalization.
4.1.2 Evaluation Metrics
We adopt a multi-dimensional evaluation framework to assess performance in sequential, conditional recommendation tasks, including:
  • Task-focused Sequential Engagement Success (TSE): A sequence-level metric that measures task completion efficiency in Task-focused dialogues, covering requirement matching, purchase-stimulating information delivery, and proactive guidance. The mathematical formula for TSE is given as follows:
    $\begin{split}\text{TSE}(k)=&\ \omega_{1}I_{1}\gamma^{N_{1}-1}+\omega_{2}I_{2}\gamma^{N_{2}-N_{1}-1}\\ &+\omega_{3}I_{3}\gamma^{N_{3}-N_{2}-1}.\end{split}$ (11)
    Here, $\omega_{1},\omega_{2},\omega_{3}$ are task weights ($\sum\omega_{i}=1$), $I_{i}$ indicate task completion, $\gamma$ is a turn-based decay factor, and $N_{i}$ denote the turns at which each task is completed. TSE directly measures the effectiveness of business decision-making by assessing whether key tasks are completed in minimal dialogue turns.
  • Reward (LM Score): Responses are scored by a large model (e.g., GPT-4) and averaged over dialogue turns.
  • BLEU(Papineni et al., 2002): Evaluates n-gram overlap with references but lacks semantic sensitivity.
  • G-Eval(Liu et al., 2023): An LLM-based metric assessing coherence, completeness, and emotional appropriateness via structured prompts.
4.1.3 Baseline Models and Comparative Experiments
We conducted two groups of comparative experiments:
  • Experiment 1: Baseline Comparison. We compare GOPO with untrained LLMs, SFT, PPO, and Memento as representative supervised and RL baselines.
  • Experiment 2: Comparison with Large Models. We benchmark GOPO against proprietary and open-source flagship models, including GPT-5.2 and Gemini-2.5 (closed-source), as well as Qwen-235B, DeepSeek-R1, and GLM-4.7 (open-source). We also evaluate GOPO implementations on Qwen-7B-Chat, Qwen3-8B, and Qwen3-14B to assess scalability across model sizes.
4.2 Main Results and Results Analysis
4.2.1 Advantages of GOPO over Mainstream Training Frameworks
Across all four datasets, GOPO consistently outperforms the untrained model and strong baselines including SFT, PPO, and Memento. Specifically, GOPO improves TSE by 25%–38% and Reward by 12%–40% over the untrained model, demonstrating the effectiveness of goal-oriented training.
Compared to SFT, GOPO achieves 10%–15% higher TSE, highlighting its advantage in sequence-conditioned, business-driven dialogue tasks beyond token-level NLG optimization. Relative to PPO, GOPO yields 7%–10% gains in TSE and 3%–20% in Reward, validating the benefit of dual-agent decomposition and business-aligned reward design over single-agent RL. Finally, GOPO outperforms Memento by 9%–12% in TSE and 3%–8% in Reward, indicating strong robustness across training paradigms.
4.2.2 Comparison with Large-Scale Frontier Models
We compare GOPO with large-scale frontier models, including Qwen-235B, DeepSeek-R1, GLM-4.7, GPT-5.2, and Gemini-2.5, and observe three key findings.
First, GOPO achieves competitive performance with significantly fewer parameters. Despite being over an order of magnitude smaller (7B/8B vs. 235B+), GOPO-Qwen-7B and GOPO-Qwen3-8B attain comparable TSE scores. On Mgshop, GOPO-Qwen3-8B reaches a TSE of 0.9375, closely matching DeepSeek-R1 (0.9387).
Second, SOP-based prompt constraints yield inconsistent gains for large models. While marginal improvements are observed for some models (e.g., Qwen-235B and GLM-4.7), others (e.g., Gemini-2.5 on Multiwoz) exhibit degraded performance, indicating the brittleness and model-dependence of prompt-level soft constraints.
Finally, GOPO demonstrates superior stability across datasets and settings. Its hierarchical dual-agent architecture and explicit constraint propagation ensure reliable enforcement of business objectives, outperforming prompt-based approaches in both robustness and consistency.
Overall, these results demonstrate that GOPO provides a more effective optimization paradigm for complex, Goal-Oriented dialogue tasks under realistic business constraints.
4.2.3 Performance Analysis under Different GOPO Parameter Settings
The training trends and experimental results of the GOPO framework under different parameter settings are shown in Figure 3 and Table 4. It can be observed that as the model scale increases, GOPO achieves improvements in TSE, Reward, and BLEU across all datasets, with GOPO-Qwen3-14B performing the best. The medium-sized GOPO-Qwen3-8B also significantly outperforms GOPO-Qwen-7B-Chat, while the performance gains vary across datasets, with more noticeable improvements on Mgshop and the TmallBrand series. Notably, despite variations in individual loss values, the overall reward converges reliably.
4.3 Ablation Study
To assess the contribution of each component in GOPO, we conduct ablation studies on Mgshop and TmallBrand-A. As shown in Table 5, we compare the full GOPO framework with a variant without the Expert Agent, as well as the corresponding untrained baseline.
Removing the Expert Agent causes substantial performance degradation, confirming its critical role in long-horizon optimization. On TmallBrand-A, TSE drops by 5.7% and Reward decreases by 13.3%. Notably, although BLEU slightly increases, TSE declines sharply, indicating that without strategic guidance the model produces conservative yet inefficient responses. This further exposes the limitation of BLEU as a surface-level metric for task-oriented dialogue.
Compared to the untrained baseline, Full GOPO achieves significant gains, improving TSE by 28.9% and 33.6%, and Reward by 10.3% and 16.0% on the two datasets. These results demonstrate that the ESNDCG-driven hierarchical RL objective effectively guides the model toward efficient task completion.
Interestingly, even without specialized training, the dual-agent architecture alone yields clear improvements over the single-agent baseline, suggesting that decoupling strategy planning from response generation better matches the structure of task-oriented dialogue.
Finally, mixed-agent experiments in Tables 6 and 7 show consistent performance drops when either agent is replaced by a non-jointly trained counterpart, confirming that GOPO’s effectiveness arises from the synergistic optimization of both agents rather than any single component.
5 Conclusion
We address the “efficiency trap” in task-oriented dialogue systems for complex business scenarios, where optimization objectives misalign with business KPIs and strategies lack flexibility and compliance. We propose GOPO, a hierarchical RL framework with a dual-agent architecture that decouples strategy planning from response execution. An Expert Agent selects high-level strategies, while a Customer Service Agent generates compliant responses under strong constraints. Combined with SOP-matching and a joint reward model aligned with business end-state metrics, GOPO shifts optimization from preference imitation to direct business value maximization. Experiments on real-world datasets show consistent gains over all baselines, including large proprietary models, on metrics such as TSE and Reward.
GOPO currently relies on a manually defined policy pool, and its reward design can be further improved with advanced credit assignment. Future work will explore automatic strategy discovery and broader domain generalization. Overall, GOPO offers a practical framework for business-oriented task dialogue and opens new directions for goal-driven multi-agent LLM optimization.